Question 1:

A lottery claims that its prize is $3 million. The first $1 million is paid immediately, and the second $1 million payment will be made in 1 year, and the third $1 million payment will be made in 2 years. What is the present value of this prize? Interest rate stays at 5% throughout the whole process.
Answer #1

ANswer

the value is $2.7 million approximately

PV=A*(P|A,i,n)

A=$1 million

PV=present value=the true price of a prize

n=years=3

i=r=interest rate =5%

PV=1*(P|A,5%,3)

=1* 2.72324802937048

=2.72324802937048

=2.723248 million

==========

with formulas

(+)-7= Ad

41-(1+0.05) 3 PV = 1* - 0.05 PV = $2.72325